Operations Report – May 9th

SOL: 4

Name of person filing report: Paolo Guardabasso

Non-nominal systems: None

Notes on non-nominal systems: Nothing to report

Generator (hours run last night):

On (last night): 8:00 PM

Off (this morning): 7:00 AM

Solar— SOC 85% at 7:20 PM

Diesel Reading – 97%

Propane Reading – 46%

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 230 gallons

Axillary to Static tank transfer – no

Gallons transferred: No

Water in GreenHab – 160 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 143406.8

Toilet tank emptied: No

Deimos rover used: no (out of order)

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 92.1h

Beginning charge:

Ending charge:

Currently charging: Yes

Opportunity rover used: No

Hours: 58.7h

Beginning charge:

Ending charge:

Currently charging: No

Curiosity rover used: No

Hours: 94.3h

Beginning charge:

Ending charge:

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? Nothing to report

CrewCar used and why, where? Off Campus

General notes and comments: Nothing to report

Summary of internet: Signal was down during the morning, then returned after lunch.

Summary of suits and radios: Some repairs were done on the suits.

Summary of Hab operations: A fantastic pesto sauce was prepared with the basil from the GreenHab

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: The damaged autoclave has been successfully used for heating (the pressure functions are unavailable). [From Director]: the binocular scope is damaged. I used it this evening and it will need some adjustment and

hopefully I will be able to fix it. I will try tomorrow morning.

Summary of RAM operations: It was depressurized to allow the EVA crew to take the drones for the experiments.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: The water heater is on, but it seems to be not functioning properly. Is there a procedure we are not doing correctly?